Grade 9Trigonometry

A pendulum bob of weight 1N is held at an angle tita from the vertical by a horizontal force of 2N. The tension is the string supporting the pendulum bob (in newton) is ???

Shamitha
8 Years agoGrade 9

1 Answer

Shamitha
8 Years ago
Thequestion clearly asks;....The tension in the string supporting the pendulum(in Newtons) bob is m×g=1NF=2N(1)T=mg/cos(tita)=1/cos(tita)(2)T=F/sin(tita)=2/sin(tita)=> T=whole root of (Tsin(tita))^2 + (Tcos(tita))^2=whole root of 2^2+1^2=whole root of 4+1=> root of 5